OpenStudy (anonymous):

Find domain, range, x and y intercept, Horizontal and vertical asymptote? How do I do it? f(x)=3^x-9

OpenStudy (anonymous):

The domain is all the possible values you can have for x. There is no value for x where we would have "problems" like dividing by 0 or having a negative under the square root. Therefore, the domain of that function is all real numbers or (-infinity, infinity). The range is all the possible values you can have for y. If you solve for x in your equation, you'll have: \[f(x)+9=3^x\] \[x = \log_{3}[f(x)+9] \] Now, what possible values can we have for f(x)? Since a log cannot have 0 or negative numbers on the inside, f(x) > -9. So, the range is (-9, infinity) There are no vertical asymptotes for exponential functions (just kind of a given if you are precalculus). The horizontal asymptote occurs on the y = -9 when we look at range.

OpenStudy (anonymous):

What about the x and y int? On my solution set i have x-int: (2,0) and y-int: (0,-8). How do you get those answers? I understand the other parts.

OpenStudy (anonymous):

Oh whoops. Forgot about the intercepts. To find the x-intercept, let y = 0 to get 0 = 3^x - 9; solve for x to get x = 2, or the point (2, 0). To find the y-intercept, let x = 0 to get f(x) = 3^0 - 9 = 1 - 9 = -8. So, the y-intercept is at the point (0, -8)